OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - y - 2z if x = -2, y = 3, and z = -2 I got 9 for my answer, but the choices are: A. -15 B. 15 C. -7 D. 7

OpenStudy (anonymous):

Did you substitute the values correctly?

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

Can you show your working? I didn't get 9

OpenStudy (anonymous):

since x= -2 y= 3 z= -2 I put that into the equation: 4(-2) - (3) - 2- (-2) from there, I multiplied, and subtracted the numbers. I got 9...

OpenStudy (anonymous):

4(-2) - 3 - 2(-2) not 4(-2) - 3 - 2 - (-2)

OpenStudy (anonymous):

I typed that wrong, in my work, that's the equation I had... I still got 9

OpenStudy (anonymous):

4(-2) - 3 - 2(-2) = -8 - 3 + 4 = -11 + 4 = -7

OpenStudy (anonymous):

i didn't combine the -8 and -3. thank you so much.